OpenStudy (anonymous):

Does sqrt{x+1} = (x+1)_{2}^{1}

OpenStudy (anonymous):

\[Does \sqrt{x+1} = (x+1)_{2}^{1}\]

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

\[\sqrt{x+1} = (x+1)^{\frac12}\]

OpenStudy (anonymous):

So does \[( (x+1)_{2}^{1})_{}^{2} = (x + 1)_{2}^{2} = x + 1 ???\]

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

Can \[(\sqrt{4x+7})_{}^{2} = (\sqrt{4x + 7})(\sqrt{4x + 7}) ??\]

OpenStudy (saifoo.khan):

Correct.

OpenStudy (anonymous):

But wouldn't multiplying those give a trinomial answer?

OpenStudy (saifoo.khan):

No

OpenStudy (saifoo.khan):

\[\sqrt{3} \times \sqrt{3} = 3\]

OpenStudy (anonymous):

There must be a rule I am missing, can only the 4x * the 4x and the 7 * the 7 Inside the radicals?

OpenStudy (saifoo.khan):

Sorry, rephrase please?

OpenStudy (anonymous):

Ill make new question

OpenStudy (saifoo.khan):

Perfect.